Lumi:

2.10 Unit 2 Test: Part 2

Lumi:

1. Renting a bicycle at a local Bike shop costs $30 for the first day and $10 for each additional day. (a) Create a chart of values and write an equation to find the total cost, C in terms of the number of days d. (b) Use the equation you created in Part (a) to find the cost of renting a bike for 5 days. (a) Days: 1, 2, 3 Cost: 30, 40, 50 Cost = 30 + 10x Where x = # of each additional day (b) 30 + 10x 30 + 10(5) 30 + 50 = 80 2. Simplify the following problem using the distributive property: 2x + 3y + 2(x - y) -3x . Show each step in your calculation. Justify each of your steps using number properties. 2x + 3y + 2(x - y) -3x 2x + 3y + 2x - 2y -3x Distributive Property 4x + 3y -2y - 3x Addition 4x + y - 3x Subtraction x + y Addition

Lumi:

If I am incorrect or you have any criticisms, please explain below.